Find the constant term that should be added to make the following expression a perfect square trinomial.
x^2 + 3x

answer choice of 9, 9/4, 9/2 or 3/2
my answer is 9/2 but am not confident it is correct

Answer by venugopalramana(3286)

Find the constant term that should be added to make the following expression a perfect square trinomial.
x^2 + 3x
TRY COMPLETING THE SQUARE FULLY USING X^2 AND X TERMS
{(X)^2+2*(3/2)*X)}IS UTILISATION OF THE GIVEN TERMS.OK NOW MAKE A PERFECT SQUARE USING FORMULA FOR (A+B)^2=A^2+2AB+B^2
{(X)^2+2*(3/2)*X)+(3/2)^2}....(3/2)^2=9/4 SHOULD BE ADDED.
answer choice of 9, 9/4, 9/2 or 3/2
my answer is 9/2 but am not confident it is correct
NO SEE ABOVE ...9/4 IS THE ANSWER.